OpenStudy (anonymous):

Help The sum of two integers is 15. Their product is 9 more than the difference of their sauares. Find the integers. If no solution explain why

OpenStudy (anonymous):

Let the integers be represented by x and y. Try to generate equations with these variables based on the information from the sentences.

OpenStudy (anonymous):

I got 7 and 8 but it doesn't fit the 2nd reason

OpenStudy (anonymous):

Oh i think 9 and 6

OpenStudy (anonymous):

The sum of two integers is 15 is the same as saying: x + y = 15 What equation can you think of that describes "The product of two integers is 9 more than the difference of their squares?"

OpenStudy (anonymous):

9+(x^2-y^2)

OpenStudy (anonymous):

Perfect! Now you have a system of two equations. What do we do next?

OpenStudy (anonymous):

Find the integers?

OpenStudy (anonymous):

That's right. I would probably use substitution here.

OpenStudy (anonymous):

But how?

OpenStudy (anonymous):

I'll substitute the number 9,6?

OpenStudy (anonymous):

From the first equation: x + y = 15 We can rewrite this as y = 15 - x For your second equation we have: x*y = 9+(x^2-(15 - x)^2) Substitute y = 15 - x into this equation. x * (15 - x) = 9 + (x^2-(15 - x)^2)